OpenStudy (junebuggie12):

Is she a skilled player for hitting the bull’s eye or is the result of hers expected?

OpenStudy (junebuggie12):

You play darts with a friend. The board is a circle with a 6-inch radius. The “Bull’s Eye” is a smaller circle in the center of the board with an 0.25-inch radius. You throw darts (small metal arrows) at the board. At the end of six weeks, she has thrown 500 darts. She hit the board every time and hit the Bull’s Eye 1 time If you throw 500 darts, how many times should you expect to hit the Bull’s Eye by random chance? Compare this result to your friend’s result. Is she a skilled player for hitting the bull’s eye or is the result of hers expected?

OpenStudy (mjdennis):

What kind of distribution are your lessons about? this is a pretty easy problem in a uniform distribution (likely to hit anywhere on the board), tough in a "normal" distribution (most likely to hit one spot...)

OpenStudy (junebuggie12):

Well since the bullseye is considerably smaller than the rest of the board I'd assume this isn't a uniform problem. I'm not sure if that's what you're asking though

OpenStudy (agent0smith):

If it's geometry, it means you need to find the area of the bullseye, then find the area of the dartboard, then divide them.

OpenStudy (junebuggie12):

oh okay! thank you

OpenStudy (junebuggie12):

I get 5.76, so what do I do with it? Is that the probability of hitting the bull's eye on average?

OpenStudy (agent0smith):

I don't know what that is; show your work.

OpenStudy (mjdennis):

Same as @agent0smith , we should look at uniform and figure what the relationship is between the the bullseye and the whole board. But I'll make a quick sidebar: Example of uniform probability: you put a plate out in the rain. a raindrop is just as likely to hit anywhere on the plate. Example of normal distribution: bell curve, like if you shoot a gun at a target. most shots will hit near the center and a few, less likely shots will hit farther out.

OpenStudy (junebuggie12):

Area of the board is 113.04 (6^2 * 3.14) and area of the bull's eye is 19.625 (0.25^2 * 3.14), 19.625/113.04 = 5.76

OpenStudy (junebuggie12):

Ah, yeah, then it's uniform probability.

OpenStudy (agent0smith):

19.625/113.04 = 5.76 Does that seem reasonable?

OpenStudy (junebuggie12):

err sorry 113.04/19.625

OpenStudy (agent0smith):

Probability is never greater than 1. Should tell you that you're doing something wrong.

OpenStudy (junebuggie12):

I figured I was dividing the wrong way around 19.625/113.04 = 0.174

OpenStudy (agent0smith):

That's the probability. Now If you throw 500 darts, how many times should you expect to hit the Bull’s Eye by random chance?

OpenStudy (junebuggie12):

0.87 times?

OpenStudy (agent0smith):

You're throwing 500 darts. On each throw, your chance of bullseye is 0.174, or 17.4% Does 0.87 times seem reasonable?

OpenStudy (junebuggie12):

oh I see

OpenStudy (junebuggie12):

So 87 times out of 500?

OpenStudy (agent0smith):

Show your work so I know if you're doing it right.

OpenStudy (junebuggie12):

Wait, 17.4% of 100 is 17.4, so is 17.4 * 5 what I'm looking for?

OpenStudy (agent0smith):

That works :)

OpenStudy (junebuggie12):

87/500 chance then

OpenStudy (agent0smith):

Remember the question asks If you throw 500 darts, how many times should you expect to hit the Bull’s Eye by random chance?

OpenStudy (agent0smith):

So how does it compare to your friend? To put it nicely, she sucks a.ss.

OpenStudy (junebuggie12):

Hahahaha yep! Thank you very much for your help.

OpenStudy (agent0smith):

Welcome.

OpenStudy (agent0smith):

@junebuggie12 0.25^2 * 3.14 that cannot be 19...

OpenStudy (agent0smith):

You left too quickly but that being wrong has made your final answer wrong.

OpenStudy (agent0smith):

Oh, it's 0.19625... not 19.625.

OpenStudy (mathmale):

Important principle: probabilities are never less than zero nor greater than one: \[0\le x \le 1\]

OpenStudy (mathmale):

So you should know that a probability result such as 19.625 must be discarded immediately. On the other hand, 0.19625 could be a probability.

OpenStudy (agent0smith):

@mathmale it wasn't her probability that was off, it was just the area of the bullseye.